\begin{thebibliography}{53}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi


\bibitem[Blackwell(1953)]{blackwell1953comparison}
D.~Blackwell.
\newblock {Equivalent Comparisons of Experiments}.
\newblock \emph{The Annals of Mathematical Statistics}, 24\penalty0
  (2):\penalty0 265 -- 272, 1953.
\newblock \doi{10.1214/aoms/1177729032}.

\bibitem[Blume-Kohout et~al.(2010)Blume-Kohout, Ng, Poulin, and
  Viola]{blume2010information}
R.~Blume-Kohout, H.~K. Ng, D.~Poulin, and L.~Viola.
\newblock Information-preserving structures: A general framework for quantum
  zero-error information.
\newblock \emph{Physical Review A}, 82\penalty0 (6):\penalty0 062306, 2010.
\newblock \doi{10.1103/PhysRevA.82.062306}.

\bibitem[Bushell(1973)]{bushell1973hilbert}
P.~J. Bushell.
\newblock Hilbert's metric and positive contraction mappings in a {Banach} space.
\newblock \emph{Archive for Rational Mechanics and Analysis}, 52:\penalty0
  330--338, 1973.
\newblock \doi{10.1007/BF00247467}.

\bibitem[Carlen and Vershynina(2018)]{carlen2018recovery}
E.~A. Carlen and A.~Vershynina.
\newblock Recovery and the data processing inequality for quasi-entropies.
\newblock \emph{IEEE Transactions on Information Theory}, 64\penalty0
  (10):\penalty0 6929--6938, 2018.
\newblock \doi{10.1109/TIT.2018.2812038}.

\bibitem[Carlen and Vershynina(2020)]{carlen2020recovery}
E.~A. Carlen and A.~Vershynina.
\newblock Recovery map stability for the data processing inequality.
\newblock \emph{Journal of Physics A: Mathematical and Theoretical},
  53\penalty0 (3):\penalty0 035204, 2020.
\newblock \doi{10.1088/1751-8121/ab5ab7}.

\bibitem[Faulkner and Hollands(2022)]{faulkner2022approximate2}
T.~Faulkner and S.~Hollands.
\newblock Approximate recoverability and relative entropy {II}: 2-positive
  channels of general von {Neumann} algebras.
\newblock \emph{Letters in Mathematical Physics}, 112\penalty0 (2):\penalty0
  26, 2022.
\newblock \doi{10.1007/s11005-022-01510-9}.

\bibitem[Faulkner et~al.(2022)Faulkner, Hollands, Swingle, and
  Wang]{faulkner2022approximate}
T.~Faulkner, S.~Hollands, B.~Swingle, and Y.~Wang.
\newblock Approximate recovery and relative entropy {I}: General von {Neumann}
  subalgebras.
\newblock \emph{Communications in Mathematical Physics}, 389\penalty0
  (1):\penalty0 349--397, 2022.
\newblock \doi{10.1007/s00220-021-04143-6}.

\bibitem[Fawzi and Renner(2015)]{fawzi2015quantum}
O.~Fawzi and R.~Renner.
\newblock Quantum conditional mutual information and approximate {Markov}
  chains.
\newblock \emph{Communications in Mathematical Physics}, 340\penalty0
  (2):\penalty0 575--611, 2015.
\newblock \doi{10.1007/s00220-015-2466-x}.

\bibitem[Frenkel(2022)]{frenkel2022integral}
P.~E. Frenkel.
\newblock Integral formula for quantum relative entropy implies data processing
  inequality.
\newblock arXiv:2208.12194, 2022.

\bibitem[Fuchs and Van De~Graaf(1999)]{fuchs1999cryptographic}
C.~A. Fuchs and J.~Van De~Graaf.
\newblock Cryptographic distinguishability measures for quantum-mechanical
  states.
\newblock \emph{IEEE Transactions on Information Theory}, 45\penalty0
  (4):\penalty0 1216--1227, 1999.
\newblock \doi{10.1109/18.761271}.

\bibitem[Gao and Wilde(2021)]{gao2021recoverability}
L.~Gao and M.~M. Wilde.
\newblock Recoverability for optimized quantum f-divergences.
\newblock \emph{Journal of Physics A: Mathematical and Theoretical},
  54\penalty0 (38):\penalty0 385302, 2021.
\newblock \doi{10.1088/1751-8121/ac1dc2}.

\bibitem[Gao et~al.(2023)Gao, Li, Marvian, and Rouz{\'e}]{gao2023sufficient}
L.~Gao, H.~Li, I.~Marvian, and C.~Rouz{\'e}.
\newblock Sufficient statistic and recoverability via {Quantum Fisher
  Information} metrics.
\newblock arXiv:2302.02341, 2023.

\bibitem[Gr{\"u}mm(1973)]{grumm1973two}
H.~Gr{\"u}mm.
\newblock Two theorems about $C_p$.
\newblock \emph{Reports on Mathematical Physics}, 4\penalty0 (3):\penalty0
  211--215, 1973.
\newblock \doi{10.1016/0034-4877(73)90026-8}.

\bibitem[Hiai(2018)]{hiai2018quantum}
F.~Hiai.
\newblock {Quantum $f$-divergences in von Neumann algebras. I. Standard
  $f$-divergences}.
\newblock \emph{Journal of Mathematical Physics}, 59\penalty0 (10):\penalty0
  102202, 2018.

\bibitem[Hiai(2021)]{hiai2021quantum}
F.~Hiai.
\newblock \emph{Quantum $f$-Divergences in von Neumann Algebras: Reversibility
  of Quantum Operations}.
\newblock Mathematical Physics Studies. Springer, Singapore, 2021.
\newblock ISBN 9789813341999.
\newblock \doi{10.1007/978-981-33-4199-9}.

\bibitem[Hiai and Mosonyi(2017)]{hiai2017different}
F.~Hiai and M.~Mosonyi.
\newblock Different quantum f-divergences and the reversibility of quantum
  operations.
\newblock \emph{Reviews in Mathematical Physics}, 29\penalty0 (07):\penalty0
  1750023, 2017.
\newblock \doi{10.1142/S0129055X17500234}.

\bibitem[Hiai et~al.(2011)Hiai, Mosonyi, Petz, and B{\'e}ny]{hiai2011quantum}
F.~Hiai, M.~Mosonyi, D.~Petz, and C.~B{\'e}ny.
\newblock Quantum f-divergences and error correction.
\newblock \emph{Reviews in Mathematical Physics}, 23\penalty0 (07):\penalty0
  691--747, 2011.
\newblock \doi{10.1142/S0129055X11004412}.

\bibitem[Holevo(1973)]{holevo1973bounds}
A.~S. Holevo.
\newblock Bounds for the quantity of information transmitted by a quantum
  communication channel.
\newblock \emph{Problemy Peredachi Informatsii}, 9\penalty0 (3):\penalty0
  3--11, 1973.

\bibitem[Jen{\v{c}}ov{\'a}(2012{\natexlab{a}})]{jencova2012comparison}
A.~Jen{\v{c}}ov{\'a}.
\newblock Comparison of quantum binary experiments.
\newblock \emph{Reports on Mathematical Physics}, 70\penalty0 (2):\penalty0
  237--249, 2012{\natexlab{a}}.
\newblock \doi{10.1016/S0034-4877(12)60043-3}.

\bibitem[Jen{\v{c}}ov{\'a}(2012{\natexlab{b}})]{jencova2012reversibility}
A.~Jen{\v{c}}ov{\'a}.
\newblock Reversibility conditions for quantum operations.
\newblock \emph{Reviews in Mathematical Physics}, 24\penalty0 (07):\penalty0
  1250016, 2012{\natexlab{b}}.
\newblock \doi{10.1142/S0129055X1250016X}.

\bibitem[Jen{\v{c}}ov{\'a}(2017)]{jencova2017preservation}
A.~Jen{\v{c}}ov{\'a}.
\newblock {Preservation of a quantum R{\'e}nyi relative entropy implies
  existence of a recovery map}.
\newblock \emph{Journal of Physics A}, 50:\penalty0 085303, 2017.
\newblock \doi{10.1088/1751-8121/aa5661}

\bibitem[Jen{\v c}ov{\'a}(2018)]{jencova2018renyi}
A.~Jen{\v c}ov{\'a}.
\newblock {R{\'e}nyi relative entropies and noncommutative $L_p$-spaces}.
\newblock \emph{Annales Henri Poincar{\'e}}, 19:\penalty0 2513--2542, 2018.
\newblock \doi{10.1007/s00023-018-0683-5}.

\bibitem[Jen{\v{c}}ov{\'a}(2021)]{jencova2021renyi}
A.~Jen{\v{c}}ov{\'a}.
\newblock {R{\'e}nyi relative entropies and noncommutative $L_p$-spaces II}.
\newblock \emph{Annales Henri Poincar{\'e}}, 22:\penalty0 3235–3254, 2021.
\newblock \doi{10.1007/s00023-021-01074-9}.

\bibitem[Jen{\v{c}}ov{\'a} and Petz(2006)]{jencova2006sufficiency}
A.~Jen{\v{c}}ov{\'a} and D.~Petz.
\newblock Sufficiency in quantum statistical inference: A survey with examples.
\newblock \emph{Infinite Dimensional Analysis, Quantum Probability and Related
  Topics}, 9\penalty0 (03):\penalty0 331--351, 2006.
\newblock \doi{10.1142/S0219025706002408}.

\bibitem[Jen{\v c}ov{\'a} and Petz(2006)]{jencova2006sufficiencyfirst}
A.~Jen{\v c}ov{\'a} and D.~Petz.
\newblock Sufficiency in quantum statistical inference.
\newblock \emph{Communications in Mathematical Physics}, 263:\penalty0
  259--276, 2006.
\newblock \doi{10.1007/s00220-005-1510-7}.

\bibitem[Jen\v{c}ov{\'a}(2010)]{jencova2010quantum}
A.~Jen\v{c}ov{\'a}.
\newblock Quantum hypothesis testing and sufficient subalgebras.
\newblock \emph{Letters in Mathematical Physics}, 93:\penalty0 15--27, 2010.
\newblock \doi{10.1007/s11005-010-0398-0}.

\bibitem[Jen\v{c}ov{\'a} and Ruskai(2010)]{jencova2010aunified}
A.~Jen\v{c}ov{\'a} and M.~B. Ruskai.
\newblock A unified treatment of convexity of relative entropy and related
  trace functions, with conditions for equality.
\newblock \emph{Reviews in Mathematical Physics}, 22:\penalty0 1099--1121,
  2010.

\bibitem[Junge et~al.(2018)Junge, Renner, Sutter, Wilde, and
  Winter]{junge2018universal}
M.~Junge, R.~Renner, D.~Sutter, M.~M. Wilde, and A.~Winter.
\newblock Universal recovery maps and approximate sufficiency of quantum
  relative entropy.
\newblock \emph{Annales Henri Poincar{\'e}}, 19\penalty0 (10):\penalty0
  2955--2978, 2018.
\newblock \doi{10.1007/s00023-018-0716-0}.

\bibitem[{Le Cam}(1964)]{lecam1964sufficiency}
L.~{Le Cam}.
\newblock Sufficiency and approximate sufficiency.
\newblock \emph{The Annals of Mathematical Statistics}, 35\penalty0
  (4):\penalty0 1419--1455, 1964.
\newblock \doi{10.1214/aoms/1177700372}.

\bibitem[Leditzky et~al.(2017)Leditzky, Rouz{\'e}, and Datta]{leditzky2017data}
F.~Leditzky, C.~Rouz{\'e}, and N.~Datta.
\newblock Data processing for the sandwiched {R}{\'e}nyi divergence: a
  condition for equality.
\newblock \emph{Letters in Mathematical Physics}, 107\penalty0 (1):\penalty0
  61--80, 2017.
\newblock \doi{10.1007/s11005-016-0896-9}.

\bibitem[Lieb and Ruskai(1973)]{lieb1973proof}
E.~H. Lieb and M.~B. Ruskai.
\newblock Proof of the strong subadditivity of quantum-mechanical entropy.
\newblock \emph{Les rencontres physiciens-math{\'e}maticiens de
  Strasbourg-RCP25}, 19:\penalty0 36--55, 1973.
\newblock \doi{10.1063/1.1666274}.

\bibitem[Lindblad(1975)]{lindblad1975completely}
G.~Lindblad.
\newblock Completely positive maps and entropy inequalities.
\newblock \emph{Communications in Mathematical Physics}, 40:\penalty0 147--151,
  1975.
\newblock \doi{10.1007/BF01609396}.

\bibitem[{\L}uczak(2014)]{luczak2014quantum}
A.~{\L}uczak.
\newblock Quantum sufficiency in the operator algebra framework.
\newblock \emph{International Journal of Theoretical Physics}, 53\penalty0
  (10):\penalty0 3423--3433, 2014.
\newblock \doi{10.1007/s10773-013-1747-4}.

\bibitem[{\L}uczak(2015)]{luczak2015onageneral}
A.~{\L}uczak.
\newblock {On a general concept of sufficiency in von Neumann algebras}.
\newblock \emph{Probab. Math. Stat.}, 35:\penalty0 313--324, 2015.

\bibitem[Mosonyi and Petz(2004)]{mosonyi2004structure}
M.~Mosonyi and D.~Petz.
\newblock Structure of sufficient quantum coarse-grainings.
\newblock \emph{Letters in Mathematical Physics}, 68:\penalty0 19--30, 2004.
\newblock \doi{10.1007/s11005-004-4072-2}.

\bibitem[M{\"u}ller-Hermes and Reeb(2017)]{mullerhermes2017monotonicity}
A.~M{\"u}ller-Hermes and D.~Reeb.
\newblock Monotonicity of the quantum relative entropy under positive maps.
\newblock \emph{Annales Henri Poincar{\'e}}, 18:\penalty0 1777--1788, 2017.
\newblock \doi{doi.org/10.1007/s00023-017-0550-9}.

\bibitem[Paulsen(2002)]{paulsen2002completely}
V.~Paulsen.
\newblock \emph{Completely Bounded Maps and Operator Algebras}.
\newblock Cambridge University Press, Cambridge, 2002.

\bibitem[Petz(1986)]{petz1986sufficient}
D.~Petz.
\newblock Sufficient subalgebras and the relative entropy of states of a von
  {Neumann} algebra.
\newblock \emph{Communications in Mathematical Physics}, 105\penalty0
  (1):\penalty0 123--131, 1986.
\newblock \doi{10.1007/BF01212345}.

\bibitem[Petz(1988)]{petz1988sufficiency}
D.~Petz.
\newblock Sufficiency of channels over von {Neumann} algebras.
\newblock \emph{The Quarterly Journal of Mathematics}, 39\penalty0
  (1):\penalty0 97--108, 1988.
\newblock \doi{10.1093/qmath/39.1.97}.

\bibitem[Pfanzagl(1974)]{pfanzagl1974acharacterization}
J.~Pfanzagl.
\newblock A characterization of sufficiency by power functions.
\newblock \emph{Metrika}, 21\penalty0 (1):\penalty0 197--199, 1974.
\newblock \doi{10.1007/BF01893900}.

\bibitem[Reeb et~al.(2011)Reeb, Kastoryano, and Wolf]{reeb2011hilbert}
D.~Reeb, M.~J. Kastoryano, and M.~M. Wolf.
\newblock Hilbert's projective metric in quantum information theory.
\newblock \emph{Journal of Mathematical Physics}, 52\penalty0 (8):\penalty0
  082201, 2011.
\newblock \doi{10.1063/1.3615729}.

\bibitem[Regula(2022)]{regula2022probabilistic}
B.~Regula.
\newblock Probabilistic transformations of quantum resources.
\newblock \emph{Physical Review Letters}, 128\penalty0 (11):\penalty0 110505,
  2022.
\newblock \doi{10.1103/PhysRevLett.128.110505}.

\bibitem[Regula et~al.(2022)Regula, Lami, and Wilde]{regula2022postselected}
B.~Regula, L.~Lami, and M.~M. Wilde.
\newblock Postselected quantum hypothesis testing, 2022.
\newblock arXiv:2209.10550

\bibitem[Ruskai(2002)]{ruskai2002inequalities}
M.~B. Ruskai.
\newblock Inequalities for quantum entropy: A review with conditions for
  equality.
\newblock \emph{Journal of Mathematical Physics}, 43\penalty0 (9):\penalty0
  4358--4375, 2002.
\newblock \doi{10.1063/1.1497701}.

\bibitem[Shirokov(2013{\natexlab{a}})]{shirokov2013reversibility}
M.~E. Shirokov.
\newblock Reversibility conditions for quantum channels and their applications.
\newblock \emph{Sbornik: Mathematics}, 204\penalty0 (8):\penalty0 1215,
  2013{\natexlab{a}}.
\newblock \doi{10.1070/SM2013v204n08ABEH004337}.

\bibitem[Shirokov(2013{\natexlab{b}})]{shirokov2013reversibilityjmp}
M.~E. Shirokov.
\newblock Reversibility of a quantum channel: general conditions and their
  applications to bosonic linear channels.
\newblock \emph{Journal of Mathematical Physics}, 54\penalty0 (11):\penalty0
  112201, 2013{\natexlab{b}}.
\newblock \doi{10.1063/1.4827436}.

\bibitem[Strasser({1985})]{strasser1985mathematical}
H.~Strasser.
\newblock \emph{{Mathematical Theory of Statistics}}.
\newblock {de Gruyter}, {Berlin, New York}, {1985}.
\newblock \doi{10.1515/9783110850826}.

\bibitem[Sutter et~al.(2016)Sutter, Tomamichel, and
  Harrow]{sutter2016strengthened}
D.~Sutter, M.~Tomamichel, and A.~W. Harrow.
\newblock Strengthened monotonicity of relative entropy via pinched {Petz}
  recovery map.
\newblock \emph{IEEE Transactions on Information Theory}, 62\penalty0
  (5):\penalty0 2907--2913, 2016.
\newblock \doi{10.1109/TIT.2016.2545680}.

\bibitem[Sutter et~al.(2017)Sutter, Berta, and
  Tomamichel]{sutter2017multivariate}
D.~Sutter, M.~Berta, and M.~Tomamichel.
\newblock Multivariate trace inequalities.
\newblock \emph{Communications in Mathematical Physics}, 352:\penalty0 37--58,
  2017.
\newblock \doi{10.1007/s00220-016-2778-5}.

\bibitem[Ticozzi and Viola(2010)]{ticozzi2010quantum}
F.~Ticozzi and L.~Viola.
\newblock Quantum information encoding, protection, and correction from
  trace-norm isometries.
\newblock \emph{Physical Review A}, 81\penalty0 (3):\penalty0 032313, 2010.
\newblock \doi{0.1103/PhysRevA.81.032313}.

\bibitem[Torgersen(1991)]{torgersen1991comparison}
E.~T\"orgersen.
\newblock \emph{Comparison of Statistical Experiments}.
\newblock Encyclopedia of Mathematics and its Applications. Cambridge
  University Press, Cambridge, 1991.
\newblock \doi{10.1017/CBO9780511666353}.

\bibitem[Uhlmann(1977)]{uhlmann1977relative}
A.~Uhlmann.
\newblock Relative entropy and the Wigner-Yanase-Dyson-Lieb concavity in an
  interpolation theory.
\newblock \emph{Commun. math. Phys}, 54:\penalty0 21--32, 1977.
\newblock \doi{https://doi.org/10.1007/BF01609834}.

\bibitem[Wilde(2015)]{wilde2015recoverability}
M.~M. Wilde.
\newblock Recoverability in quantum information theory.
\newblock \emph{Proceedings of the Royal Society A: Mathematical, Physical and
  Engineering Sciences}, 471\penalty0 (2182):\penalty0 20150338, 2015.
\newblock \doi{10.1098/rspa.2015.0338}.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{hyperref}
\usepackage{amsmath, amssymb, amsthm}
\usepackage[sort&compress,numbers]{natbib}
\usepackage{doi}
\usepackage[margin=0.8in]{geometry}

\usepackage{amsmath, amssymb, amsthm, mathtools}

\newtheorem{lemma}{Lemma}
\newtheorem{theorem}{Theorem}
\newtheorem{coro}{Corollary}

\theoremstyle{definition}
\newtheorem{defi}{Definition}

\theoremstyle{remark}
\newtheorem{remark}{Remark}

\def\Ha{\mathcal H}
\def\Te{\mathcal T}
\def\Ka{\mathcal K}
\def \Tr{\mathrm{Tr}\,}
\def\Se {\mathcal S}
\def\supp{\mathrm{supp}}
\def\<{\langle\,}
\def\>{\,\rangle}

\title{Recoverability of quantum channels via hypothesis testing}
\author{Anna Jen\v cov\'a\footnote{Mathematical Institute, Slovak Academy of Sciences, \v
Stef\'anikova 49, 814 73 Bratislava, Slovakia, jenca@mat.savba.sk}}
\date{}
\begin{document}

\maketitle

\abstract{A quantum channel is sufficient with respect to a set of input states if it can
be reversed on this set. In the approximate version, the input states can be recovered
within an error bounded by the decrease of the relative entropy under the channel.  
Using a new integral representation of the relative entropy in
[arxiv:2208.12194], we present an easy proof of a characterization  of sufficient quantum channels
and recoverability by preservation of optimal success probabilities in hypothesis testing
problems, equivalently, by preservation of $L_1$-distance.}

\section{Introduction}

One of the fundamental properties of quantum relative entropy is monotonicity under
quantum channels, or the data processing inequality (DPI):
\begin{equation}\label{eq:dpi}
D(\Phi(\rho)\|\Phi(\sigma))\le D(\rho\|\sigma)
\end{equation}
for any pair of quantum states $\rho,\sigma$ and any completely positive  trace preserving
map, \cite{lindblad1975completely,uhlmann1977relative}. The DPI implies other important quantum
entropic inequalities, such as the Holevo bound \cite{holevo1973bounds}, strong subadditivity of
von Neumann entropy (SSA) \cite{lieb1973proof} or the  joint convexity of relative
entropy. In fact, SSA, joint convexity and DPI are all equivalent, see
\cite{ruskai2002inequalities} and the proof of DPI in 
\cite{lindblad1975completely} is based on the SSA.

The question when the data processing inequality becomes an equality for a completely
positive map and a pair of states was first
answered by Petz \cite{petz1986sufficient,petz1988sufficiency}, who proved that, provided the
relative entropy $D(\rho\|\sigma)$ is finite, equality occurs if and only if the two
states can be fully recovered. This means that there exists a channel $\Psi$ such that 
$\Psi\circ\Phi(\rho)=\rho$ and $\Psi\circ\Phi(\sigma)=\sigma$. In this case, we say that
the channel $\Phi$ is sufficient with respect to the pair of states
$\{\rho,\sigma\}$, in analogy with the classical notion of a statistic sufficient with
respect to a family of probability distributions. Moreover, Petz proved that there exists
a universal recovery channel $\Phi_\sigma$, such that
$\Phi_\sigma\circ\Phi(\sigma)=\sigma$ and we have $\Phi_\sigma\circ \Phi(\rho)=\rho$ if
and only if the channel $\Phi$ is sufficient with respect to $\{\rho,\sigma\}$. 

Sufficiency, or sometimes called reversibility, of channels was studied in a number of
subsequent works and several characterizations and applications were found,
\cite{mosonyi2004structure,jencova2006sufficiencyfirst, jencova2006sufficiency,
hiai2011quantum, jencova2012reversibility, 
shirokov2013reversibilityjmp, luczak2014quantum}. In particular, equality
conditions for various forms of DPI were studied, e.g.  \cite{ruskai2002inequalities,
jencova2010aunified, leditzky2017data}, and their relation to sufficiency were examined for other information
theoretic or statistical quantities, such as  different versions of quantum
$f$-divergences \cite{hiai2017different, hiai2021quantum}, quantum R\'enyi relative
entropies \cite{jencova2017preservation,jencova2018renyi,jencova2021renyi,hiai2021quantum}, Holevo quantity \cite{shirokov2013reversibility}, Fisher information
and $L_1$-distance \cite{jencova2012reversibility}.

An approximate version of sufficiency, called (approximate) recoverability is a much
stronger result stating that if the decrease in the relative entropy is small, there
exists a channel that recovers $\sigma$  perfectly while $\rho$ is recovered up to a small error. First result of this form was proved in
the work of Fawzi and Renner \cite{fawzi2015quantum}, who considered approximate quantum
Markov chains. This was soon extended to more general channels
\cite{wilde2015recoverability,sutter2016strengthened, sutter2017multivariate,
junge2018universal} and a variety of quantities such as $f$-divergences
\cite{carlen2018recovery,carlen2020recovery}, optimized $f$-divergences
\cite{gao2021recoverability} and Fisher information \cite{gao2023sufficient}. An important
result in this context is existence of an universal recovery channel $\Phi^u_\sigma$ depending only on the
state $\sigma$ such that \cite{junge2018universal}
\[
D(\rho\|\sigma)-D(\Phi(\rho)\|\Phi(\sigma))\ge -2\log F(\rho,
\Phi^u_\sigma\circ\Phi(\rho))\ge \|\rho-\Phi^u_\sigma\circ\Phi(\rho)\|_1^2,
\]
here $\|\cdot\|_1$ denotes the trace norm and
$F(\rho,\sigma)=\|\rho^{1/2}\sigma^{1/2}\|_1$ is the fidelity. See also
\cite{gao2021recoverability,faulkner2022approximate,faulkner2022approximate2} for the respective results for
normal states of von Neumann algebras.

In the simplest setting of quantum hypothesis testing, the null hypothesis $H_0=\sigma$ is tested
against the alternative $H_1=\rho$. The tests are represented by operators $0\le M\le I$,
with the interpretation that $\Tr[\omega M]$ is the probability of rejecting the
hypothesis if the true state is $\omega$.
For the  test represented by $M$, the  Bayes error probability for $\lambda\in [0,1]$ is
expressed as
\[
P_e(\lambda,\rho,\sigma,M)=\lambda\Tr[\sigma M]+(1-\lambda)\Tr[\rho(I-M)]
\]
and the test is Bayes optimal for $\lambda$ if this error probability is minimal over all
possible tests. It is quite clear that if we replace the states by $\Phi(\sigma)$ and
$\Phi(\rho)$, the achievable error probabilities cannot be decreased. It is a natural
question when the optimal error probabilities are preserved under $\Phi$, which is
equivalent to  preservation of the $L_1$-norm:
\begin{equation}\label{eq:l1}
\|\rho-s\sigma\|_1=\|\Phi(\rho)-s\Phi(\sigma)\|_1,\qquad \forall s.
\end{equation}
In classical
statistics, the theorem of Pfanzagl \cite{pfanzagl1974acharacterization,
strasser1985mathematical} states that if the achievable  error probabilities
for a pair of probability measures
$\{P_0,P_1\}$ do not increase after transformation by a statistic $T$, then $T$ must be sufficient with respect to
$\{P_0,P_1\}$. The corresponding result for quantum channels was investigated  in
\cite{jencova2010quantum, jencova2012reversibility},  and in \cite{luczak2015onageneral},
where more general risk functions for decision problems were considered. The equivalent
question of preservation of the $L_1$-distance, with applications to error correction,
was studied in \cite{blume2010information, ticozzi2010quantum}. In all these
works, additional conditions were needed, such as the equalities have to be assumed either for
larger sets of states with a special structure, or for any number of copies of $\rho$ and $\sigma$. 
The case when $\rho$ and $\sigma$ commute, or the channel $\Phi$ has commutative range,
was solved in \cite{jencova2012reversibility}.

Many of the results on recoverability of channels rely on an integral representation of
the relative entropies or  other  quantities in question such as $f$-divergences. These
 formulas are based on integral representation of operator convex functions.
Recently, a new integral formula for the relative entropy of positive semidefinite matrices
matrices was proved in
\cite{frenkel2022integral}. This formula can be easily extended to infinite dimensional
Hilbert spaces and rewritten in terms of the optimal
Bayes error probabilities. We use this formula for simple proofs of a
characterization of recoverability of quantum channels by preserving hypothesis testing
error probabilities, or equivalently $L_1$-distances, without any additional assumptions needed in the
previous works.

\section{Preliminaries}

Throughout this paper, $\Ha$ is a 
Hilbert space and we denote by $\Te(\Ha)$ the set of operators with finite trace and by $\Se(\Ha)$ the set of states (density operators) on $\Ha$,
that is, positive operators of trace 1. For a self-adjoint operator $A\in B(\Ha)$, $A_\pm$
denotes the positive/negative part of $A$ and for $A\ge 0$, we denote the projection
onto the support of $A$ by $\supp(A)$. The $L_1$-distance in $\Te(\Ha)$ is defined as
\[
\|S\|_1:=\sup_{\|A\|\le 1} \Tr[AS]=\Tr|S|,\qquad S\in \Te(\Ha).
\]
If $S\in \Te(\Ha)$ is self-adjoint, then we have
\[
\Tr[S_+]=\sup_{0\le M\le I} \Tr[MS],\quad \Tr[S_-]=-\inf_{0\le M\le I} \Tr[MS]
\]
and 
$\Tr[S_\pm]=\tfrac12(\|S\|_1\pm\Tr[S])$.

A quantum channel $\Phi$ is a completely
positive trace preserving map  $\Te(\Ha)\to \Te(\Ka)$.  The adjoint of $\Phi$ is the map $\Phi^*:B(\Ka)\to B(\Ha)$, defined by
\[
\Tr[\Phi^*(A)\rho]=\Tr[A\Phi(\rho)],\qquad A\in B(\Ha),\ \rho\in \Se(\Ha).
\]
It is easily seen that $\Phi^*$ is completely positive and unital. 

For positive operators  $\rho,\sigma \in \Te(\Ha)$, the quantum relative entropy is defined as
\[
D(\rho\|\sigma)=\begin{dcases} \Tr[\rho(\log(\rho)-\log(\sigma))], &  \supp(\rho)\le \supp(\sigma)\\
\infty, & \text{otherwise}.
\end{dcases}
\]
Relative entropy satisfies the data processing inequality \eqref{eq:dpi} which
holds for any pair of states $\rho,\sigma$ and any quantum channel $\Phi: \Te(\Ha)\to
\Te(\Ka)$.

\subsection{Quantum hypothesis testing and $L_1$-distance}

In the problem of hypothesis testing, the task is to test the hypothesis $H_0=\sigma$ against the
alternative $H_1=\rho$. Any test is represented by  an effect on $\Ha$, that is, an
operator $0\le M\le I$, corresponding to rejecting the hypothesis. For a test $M$, the error probabilities
are
\[
\alpha(M)=\Tr[\sigma M],\qquad \beta(M)= \Tr[\rho(I-M)].
\]
For $\lambda\in (0,1)$, we define the Bayes optimal test as the minimizer of 
\[
P_e(\lambda,\sigma,\rho,M):=\lambda\alpha(M)+(1-\lambda)\beta(M)=(1-\lambda)(1-\Tr[(\rho-s\sigma)M]),\qquad
s=\frac{\lambda}{1-\lambda}.
\]
The proof of the following description of the Bayes optimal tests can be found in
\cite{jencova2010quantum}.

\begin{lemma}[Quantum Neyman-Pearson] \label{lemma:qnm} Let $\rho,\sigma$ be states,
$\lambda \in [0,1]$ and put $s=\frac{\lambda}{1-\lambda}$. A test $M$ is a Bayes optimal
test for $\lambda,\sigma,\rho$ if and only if
\[
M=P_{s,+}+X_s,\qquad 0\le X_s\le P_{s,0},
\]
where $P_{s,\pm}=\supp((\rho-s\sigma)_\pm)$ and
$P_{s,0}=I-P_{s,+}-P_{s,-}$. The optimal error probability is then
\begin{align*}
P_e(\lambda,\sigma,\rho):=\max_M
P_e(\lambda,\sigma,\rho,M)&=(1-\lambda)(1-\Tr[(\rho-s\sigma)_+])\\
&=(1-\lambda)(s-\Tr[(\rho-s\sigma)_-])\\
&=\frac12(1-(1-\lambda)\|\rho-s\sigma\|_1).
\end{align*}

\end{lemma}

It is easily seen that the error probabilities and the related quantities in the above lemma
are monotone under channels, in particular,
\begin{align*}
P_e(\lambda,\Phi(\sigma),\Phi(\rho))&\ge P_e(\lambda,\sigma,\rho),\\
\|\Phi(\rho)-s\Phi(\sigma)\|_1&\le \|\rho-s\sigma\|_1,\\
\Tr[(\Phi(\rho)-s\Phi(\sigma))_-]&\le \Tr[(\rho-s\sigma)_-].
\end{align*}
In fact, monotonicity holds if $\Phi$ is a positive trace preserving map, so complete
positivity is not needed.

\subsection{Integral formula for the relative entropy}

The following new integral representation of the relative entropy was proved by Frenkel  in
\cite{frenkel2022integral}, in the case $\dim(\Ha)<\infty$.

\begin{theorem}\label{thm:frenkel} Let $\rho,\sigma$ be positive operators in $\Te(\Ha)$. Then
\[
D(\rho\|\sigma)=\Tr[\rho-\sigma]+\int_{-\infty}^\infty
\frac{dt}{|t|(1-t)^2}\Tr[((1-t)\rho+t\sigma)_-].
\]
\end{theorem}

\begin{proof} By \cite[Theorem 6]{frenkel2022integral}, the equality holds if
$\dim(\Ha)<\infty$. We will now prove that it can be extended to the case when $\dim(\Ha)=\infty$.
Assume first that $\supp(\rho)\le \supp(\sigma)$, so that we may assume that $\sigma$ is
faithful and $\Ha$ is separable by restriction to the support of $\sigma$. We will use a standard limiting
argument to extend the finite dimensional result to the separable case.

Let $P_n$ be an increasing sequence of finite rank projections such that $\vee_n P_n=I$.
Put $\rho_n=P_n\rho P_n$, $\sigma_n=P_n\sigma P_n$. Then restricted to the finite
dimensional space $P_n\Ha$, $\rho_n$ and $\sigma_n$ are positive semidefinite operators
with $\supp(\rho_n)\le \supp(\sigma_n)$.   Moreover, $\lim_n
\Tr[\rho_n]=\Tr[\rho]$, $\lim_n\Tr[\sigma_n]=\Tr[\sigma]$ and by \cite[Theorem 4.5]{hiai2018quantum} we have
$D(\rho\|\sigma)=\lim_n D(\rho_n\|\sigma_n)$.

For $t\in \mathbb R$ and $n\in \mathbb N$, put
\[
f_n(t):=\Tr[((1-t)\rho_n-t\sigma_n)_-], \qquad f(t):=\Tr[((1-t)\rho-t\sigma)_-].
\]
Then
\begin{align*}
f_n(t)&=\Tr[(t\sigma_n-(1-t)\rho_n)_+]=\sup_{0\le M_n\le P_n}
\Tr[M_n(t\sigma-(1-t)\rho)]\\ &\le \sup_{0\le M_{n+1}\le
P_{n+1}}\Tr[M_{n+1}(t\sigma-(1-t)\rho)]=\Tr[((1-t)\rho_{n+1}-t\sigma_{n+1})_-]=f_{n+1}(t),
\end{align*}
where the inequality follows from $0\le M_n\le P_n\le P_{n+1}$. Furthermore,  since
$P_n\to I$ in the strong operator topology, we have using \cite[Theorem 1]{grumm1973two} that 
 \[
\|P_n((1-t)\rho-t\sigma)P_n\|_1\to \|(1-t)\rho-t\sigma\|_1.
\]
It follows that
\begin{align*}
f_n(t)&=\tfrac12(\|(1-t)\rho_n-t\sigma_n\|_1-\Tr[(1-t)\rho_n-t\sigma_n])\to\tfrac12
(\|(1-t)\rho-t\sigma\|_1-\Tr[(1-t)\rho-t\sigma))\\
&=f(t).
\end{align*}
Hence $f_n$ is an increasing sequence of positive integrable functions converging
pointwise to $f$.  Since the integral formula holds in finite dimensions, we see using the Lebesgue monotone
convergence theorem that
\begin{align*}
D(\rho\|\sigma)&=\lim_n D(\rho_n\|\sigma_n)=\lim_n
\left(\Tr[\rho_n-\sigma_n]+\int_{-\infty}^\infty\frac{dt}{|t|(1-t)^2}f_n(t)\right)\\
&=\Tr[\rho-\sigma]+\int_{-\infty}^\infty
\frac{dt}{|t|(1-t)^2}f(t).
\end{align*}

If $\supp(\rho)\not\le\supp(\sigma)$, then there is some projection $Q$ such that $\Tr[\sigma
Q]=0$ and $c:=\Tr[\rho Q]>0$. Then for any $t>1$ we have 
\[
\Tr[((1-t)\rho-t\sigma)_-]\ge \Tr[Q(t\sigma-(1-t)\rho)]=(t-1)c
\]
and hence
\[
\int_{-\infty}^{\infty}\frac{dt}{|t|(t-1)^2}\Tr[((1-t)\rho-t\sigma)_-]\ge
c\int_1^\infty\frac{dt}{t(t-1)}=\infty.
\]
In this case we also have $D(\rho\|\sigma)=\infty$ by definition.

\end{proof}

The integral  formula leads to an easy proof of the fact that DPI for the relative entropy holds
for all positive trace preserving maps. This fact was first
proved in \cite{mullerhermes2017monotonicity}, using interpolation techniques.

\begin{coro}\label{coro:integral} Let $\rho,\sigma\in \Se(\Ha)$. Then for any $\lambda,\mu\ge 0$ such that 
$\mu\sigma\le \rho\le \lambda \sigma$, we have 
\[
D(\rho\|\sigma)=\int_\mu^\lambda \frac{ds}{s}\Tr[(\rho-s\sigma)_-]+ \log(\lambda)+1-\lambda.
\]

\end{coro}

\begin{proof} Since $((1-t)\rho+t\sigma)_-=0$
for $t\in [0,1]$, the integral splits into two parts, integrating over $t\le 0$ and $t\ge
1$. For the first integral, since $1-t>0$, we have
$((1-t)\rho+t\sigma)_-=(1-t)(\rho-\frac{t}{t-1}\sigma)_-$ and 
\begin{align*}
\int_{-\infty}^0\frac{dt}{-t(1-t)^2}\Tr[(1-t)\rho+t\sigma]_-&=\int_{-\infty}^0
\frac{dt}{t(t-1)}\Tr[(\rho-\frac{t}{t-1}\sigma)_-]=\int_0^1\frac{ds}{s}\Tr[(\rho-s\sigma)_-]\\
&=\int_\mu^1\frac{ds}{s}\Tr[(\rho-s\sigma)_-].
\end{align*}

For $t\ge 1$, we use
$((1-t)\rho+t\sigma)_-=((t-1)\rho-t\sigma)_+=(t-1)(\rho-\frac{t}{t-1}\sigma)_+$ and
inserting into the integral, we obtain
\[
\int_1^\infty\frac{dt}{t(t-1)^2}\Tr[((1-t)\rho+t\sigma)_-]=\int_1^\infty
\frac{ds}{s}\Tr[(\rho-s\sigma)_+]=\int_1^\lambda \frac{ds}{s}\Tr[(\rho-s\sigma)_+].
\]
The proof is finished by using  the equality $\Tr[(\rho-s\sigma)_+]=1-s +\Tr[(\rho-s\sigma)_-]$.
\end{proof}

\begin{remark}\label{rem:hilbert}
The smallest value of $\lambda$ in the above expression is related to the quantum max-relative entropy
defined as\[
D_{\max}(\rho\|\sigma):=\log\min\{\lambda,\ \rho\le \lambda\sigma\}.
\]
Similarly, the largest value of $\mu$ is $e^{-D_{\max}(\sigma\|\rho)}$. An important
related quantity is the Hilbert projective metric \cite{bushell1973hilbert}
\[
D_{\Omega}(\rho\|\sigma):=D_{\max}(\rho\|\sigma)+D_{\max}(\sigma\|\rho).
\]
See \cite{reeb2011hilbert,regula2022probabilistic, regula2022postselected} for more details and interpretations in the context of
quantum information theory. Note also that we may always put $\mu=0$ and if $\dim(\Ha)<\infty$, then the conditions that
$\rho\le \lambda\sigma$ for some $\lambda>0$ is equivalent to $\supp(\rho)\le
\supp(\sigma)$, so it holds whenever $D(\rho\|\sigma)$ is finite.
In infinite dimensions, this condition is much more restrictive. 

\end{remark}

\subsection{Sufficiency and recoverability for quantum channels}

The following definition first appeared in \cite{petz1988sufficiency} and can be seen as a
quantum generalization of the classical notion of a sufficient statistic.

\begin{defi}\label{def:sufficiency}
We say that a channel $\Phi:B(\Ha)\to B(\Ka)$ is sufficient with respect to a set of
states $\Se\subseteq \Se(\Ha)$ if there exists a channel $\Psi:B(\Ka)\to B(\Ha)$ such that
\[
\Psi\circ \Phi(\rho)=\rho,\qquad  \forall \rho\in \Se.
\]
\end{defi}

For a state $\sigma\in \Se(\Ha)$, we define an inner product $\<\cdot,\cdot\>_\sigma$ in
$B(\supp(\sigma))$ by
\[
\<A,B\>_\sigma:=\Tr[A^*\sigma^{1/2}B\sigma^{1/2}],\qquad A,B\in B(\supp(\sigma)).
\]
It was proved in \cite{petz1988sufficiency} that the (unique) linear map  $\Phi_\sigma: \Te(\supp(\Phi(\sigma)))\to
\Te(\supp(\sigma))$ determined by
\[
\<\Phi^*(B),A\>_\sigma=\<B,\Phi_\sigma^*(A)\>_{\Phi(\sigma)},\qquad A\in B(\supp(\sigma)),
B\in B(\supp(\Phi(\sigma)))
\]
is a channel, called the Petz dual of $\Phi$ with respect to $\sigma$ (or the Petz
recovery map). Note that we always have $\Phi_\sigma\circ
\Phi(\sigma)=\sigma$ and as it was further proved in \cite{petz1988sufficiency}, if both
$\sigma$ and $\Phi(\sigma)$ are faithful, then $\Phi$ is sufficient with respect to $\Se$
if and only if $\Phi_\sigma\circ \Phi(\rho)=\rho$ for all $\rho\in \Se$, so that
$\Phi_\sigma$ is a universal recovery channel.
\begin{remark} If $\dim(\Ha)<\infty$, we obtain the familiar form of the Petz recovery
channel:
\[
\Phi_\sigma(\cdot)=
\sigma^{1/2}\Phi^*(\Phi(\sigma)^{-1/2}\cdot \Phi(\sigma)^{-1/2})\sigma^{1/2}.
\]
\end{remark}

We also define 
\[
\Phi_{\sigma,t}(\cdot)= \sigma^{-it}\Phi_\sigma(\Phi(\sigma)^{it}\cdot
\Phi(\sigma)^{-it})\sigma^{it},\qquad t\in \mathbb R
\]
and
\[
\Phi_{\sigma,\mu}(\cdot)=\int_{-\infty}^\infty\Phi_{\sigma,t}(\cdot)d\mu(t),
\]
for a probability measure $\mu$ on $\mathbb R$. Clearly, all these maps are channels $\Te(\supp(\Phi(\sigma)))\to
\Te(\supp(\sigma))$ that recover the state $\sigma$.

%

\begin{theorem}\label{thm:petz}  Assume that $\rho,\sigma \in \Se(\Ha)$ are such that 
$D(\rho\|\sigma)<\infty$. Then the following are equivalent.
\begin{enumerate}
\item[(i)] $\Phi$ is sufficient with respect to $\{\rho,\sigma\}$;
\item[(ii)] $D(\Phi(\rho)\|\Phi(\sigma))=D(\rho\|\sigma)$;
\item[(iii)] $\Phi_{\sigma,t}\circ \Phi(\rho)=\rho$, for some $t\in
\mathbb R$;
\item[(iv)] $\Phi_{\sigma,t}\circ \Phi(\rho)=\rho$, for  all $t\in
\mathbb R$;
\item[(v)] $\Phi_{\sigma,\mu}\circ \Phi(\rho)=\rho$ for some probability measure $\mu$.

\end{enumerate}

\end{theorem}

\begin{proof} In finite dimensions the proof follows from  \cite[Theorem 3.3]{wilde2015recoverability}. 
The proof in the general case will be given in the Appendix.

\end{proof}

The following is an approximate version of sufficiency of channels, called recoverability
of $\Phi$.

\begin{theorem}\label{thm:recoverability} \cite{junge2018universal} Let $\sigma\in
\Se(\Ha)$. Then for any channel $\Phi: \Te(\Ha)\to \Te(\Ka)$ there exists a channel
$\Phi^u_\sigma: \Te(\Ka)\to \Te(\Ha)$ such that 
$\Phi^u_\sigma\circ\Phi(\sigma)=\sigma$ and for any $\rho\in \Se(\Ha)$ we have
\[
D(\rho\|\sigma)\ge D(\Phi(\rho)\|\Phi(\sigma))-2\log F(\rho,\Phi^u_\sigma\circ
\Phi(\rho))\ge D(\Phi(\rho)\|\Phi(\sigma))+ \frac14
\|\rho-\Phi^u_\sigma\circ\Phi(\rho)\|_1^2.
\]
\end{theorem}

In the above theorem, $F(\rho_0,\rho_1)$ is the fidelity
\[
F(\rho_0,\rho_1)=\|\rho_0^{1/2}\rho_1^{1/2}\|_1.
\]
The second inequality in Theorem \ref{thm:recoverability} is
obtained using the inequality $-\log(x)\ge 1-x$ for $x\in
(0,1)$ and the Fuchs-van de Graaf inequality, \cite{fuchs1999cryptographic}.

The universal recovery channel $\Phi^u_\sigma$ can be chosen as
\begin{equation}\label{eq:universal}
\Phi^u_\sigma(\cdot)=\Phi_{\sigma,\beta_0}(P\cdot P)+\Tr[(I-P)\cdot ],
\end{equation}
here $P=\supp(\Phi(\sigma))$ and $\beta_0$ is the probability density
function
\[
\beta_0(t)=\frac{\pi}{\cosh(2\pi t)+1}.
\]
Note that if $\supp(\rho)\le \supp(\sigma)$, then $\supp(\Phi(\rho))\le
\supp(\Phi(\sigma))$, so that
$\Phi^u_\sigma(\Phi(\rho))=\Phi_{\sigma,\beta_0}(\Phi(\rho))$ and the statement in this
case follows by \cite[Theorem 2.1]{junge2018universal}. If
$\supp(\rho)\not\le\supp(\sigma)$, then $D(\rho\|\sigma)=\infty$ and the inequality holds
trivially.

\section{Sufficiency and recoverability by hypothesis testing}

The characterization in Theorem \ref{thm:petz} and the integral formula in Corollary
\ref{coro:integral} now give an easy proof of characterization of sufficiency and
recoverability by quantities related to hypothesis testing. Note that here we do not have
to make any further assumptions about the states.

\begin{theorem}\label{thm:sufficiency} Let $\Phi:\Te(\Ha)\to \Te(\Ka)$ be a channel and let
$\rho,\sigma\in \Se(\Ha)$.
Then the following are
equivalent.
\begin{enumerate}
\item[(i)] $P_e(\lambda, \Phi(\rho),\Phi(\sigma))= P_e(\lambda,\rho,\sigma)$, for all  $\lambda\in
[0,1]$;
\item[(ii)]  $\|\Phi(\rho)-s\Phi(\sigma)\|_1= \|\rho-s\sigma\|_1$, for all
$s\ge 0$;
\item[(iii)]  $\Tr[(\Phi(\rho)-s\Phi(\sigma))_{+}]= \Tr[(\rho-s\sigma)_{+}]$, for all $s\ge 0$;

\item[(iv)]  $\Tr[(\Phi(\rho)-s\Phi(\sigma))_{-}]= \Tr[(\rho-s\sigma)_{-}]$, for all
$s\ge 0$;
\item[(v)] $\Phi$ is sufficient with respect to $\{\rho,\sigma\}$.

\end{enumerate}

\end{theorem}

\begin{proof} The equivalences between (i)-(iv) are clear from Lemma \ref{lemma:qnm}. 
 Assume that (iv) holds. Suppose first that $\rho\le \lambda\sigma$ for some
$\lambda>0$. Then also $\Phi(\rho)\le
\lambda\Phi(\sigma)$ and we have by Corollary \ref{coro:integral}
\begin{align*}
D(\Phi(\rho)\|\Phi(\sigma))&=\int_0^\lambda
\frac{ds}{s}\Tr[(\Phi(\rho)-s\Phi(\sigma))_-]+\log(\lambda)+1-\lambda\\
&=
\int_0^\lambda
\frac{ds}{s}\Tr[(\rho-s\sigma)_-]+\log(\lambda)+1-\lambda=D(\rho\|\sigma).
\end{align*}
 By Theorem
\ref{thm:petz}, this implies (v). In the general case, let
$\sigma_0=\frac12(\rho+\sigma)$, then $\rho\le 2\sigma_0$ and it is easily
seen that the equality (ii) implies a similar equality with $\sigma$ replaced by
$\sigma_0$. It follows that $\Phi$ is sufficient with respect to $\{\rho,\sigma_0\}$,
which implies (v).  The implication (v) $\implies$ (ii) follows from
monotonicity of the $L_1$-distance.

\end{proof}

We are now interested in a  similar result for recoverability. Assume first that there is a channel
$\Lambda: \Te(\Ka)\to \Te(\Ha)$ such that $\Lambda\circ \Phi(\sigma)=\sigma$ and $\|\Lambda\circ
\Phi(\rho)-\rho\|_1\le \epsilon$. We then have
\begin{equation}\label{eq:epsilon_norm}
\|\rho-s\sigma\|_1=\|\rho-\Lambda\circ\Phi(\rho)+\Lambda\circ\Phi(\rho)-s\Lambda\circ
\Phi(\sigma)\|_1\le \|\Phi(\rho)-s\Phi(\sigma)\|_1+\epsilon.
\end{equation}
Using Lemma \ref{lemma:qnm}, we see that the resulting inequality in \eqref{eq:epsilon_norm} is equivalent to any of
the following inequalities
\begin{align}
\Tr[(\Phi(\rho)-s\Phi(\sigma))_+]&\ge \Tr[(\rho-s\sigma)_+]-\frac{\epsilon}2,\qquad s\ge 0\\
\Tr[(\Phi(\rho)-s\Phi(\sigma))_-]&\ge
\Tr[(\rho-s\sigma)_-]-\frac{\epsilon}2\label{eq:epsilon_minus},\qquad s\ge 0\\
P_e(\lambda,\Phi(\sigma),\Phi(\rho))&\le
P_e(\lambda,\sigma,\rho)+\frac{1-\lambda}2\epsilon,\qquad \lambda\in [0,1].
\end{align}

The following result gives the converse statement. Note that here we will need the
assumption that the Hilbert projective metric $D_\Omega(\rho\|\sigma)$ is finite,
equivalently, that $\mu\sigma\le \rho\le
\lambda\sigma$ for some $\mu,\lambda>0$ (see Remark \ref{rem:hilbert}), to get a nontrivial
result.

\begin{theorem}\label{thm:recoverability_e}  Let $\rho,\sigma\in \Se(\Ha)$ and let
$\Phi:\Te(\Ha)\to \Te(\Ka)$ be a quantum  channel. If 
\[
\|\Phi(\rho)-s\Phi(\sigma)\|_1\ge \|\rho-s\sigma\|_1-\epsilon,\qquad \forall s\ge 0
\]
holds for some $\epsilon\ge 0$, then there exists a channel $\Lambda:\Te(\Ka)\to \Te(\Ha)$ such
that $\Lambda\circ \Phi(\sigma)=\sigma$ and 
\[
\|\Lambda\circ \Phi(\rho)-\rho\|_1\le
\sqrt{2\epsilon}D_\Omega(\rho\|\sigma)^{1/2}.
\]
Moreover, we may take $\Lambda =\Phi^u_\sigma$ as in \eqref{eq:universal}.

\end{theorem}

\begin{proof} The statement is trivial if $D_\Omega(\rho\|\sigma)=\infty$, so assume that 
 $\mu\sigma\le \rho\le \lambda\sigma$ for $\mu,\lambda>0$,
 $\mu=e^{-D_{\max}(\sigma\|\rho)}$ and $\lambda=e^{D_{\max}(\rho\|\sigma)}$. Then also $\mu \Phi(\sigma)\le
\Phi(\rho)\le \lambda\Phi(\sigma)$. By the assumptions, inequality \eqref{eq:epsilon_minus} 
holds. Using  Corollary \ref{coro:integral}, we get
\begin{align*}
D(\rho\|\sigma)-D(\Phi(\rho)\|\Phi(\sigma))&=\int_\mu^\lambda\frac{ds}{s}\bigl(\Tr[(\rho-s\sigma)_-]-\Tr[(\Phi(\rho)-s\Phi(\sigma))_-]\bigr)\\
&\le \frac{\epsilon}2
\int_\mu^\lambda\frac{1}{s}ds=\frac{\epsilon}2\bigl(\log(\lambda)-\log(\mu)\bigr)=\frac{\epsilon}2D_\Omega(\rho\|\sigma).
\end{align*}
The statement now
follows by Theorem \ref{thm:recoverability}.

\end{proof}

\begin{remark} The recoverability result can be also formulated in the setting of
comparison of statistical experiments, which is an extension of the classical theory of
Blackwell \cite{blackwell1953comparison}, T\"orgersen \cite{torgersen1991comparison} and
Le Cam \cite{lecam1964sufficiency}. A
(quantum)
statistical experiment is any parametrized family of (quantum) states. For two experiments
$\mathcal E$ and $\mathcal E_0$ with the same parameter set (not necessarily living on the
same Hilbert space), 
we say that $\mathcal E_0$ is $(2,\epsilon)$-deficient with respect to $\mathcal E$ if the error
probabilities of testing problems involving elements of $\mathcal E_0$ are up to $\epsilon$ not
worse than those of corresponding testing problems for $\mathcal E$. See  \cite{jencova2012comparison} for a more precise definition. In particular, for
$\mathcal E_0=\{\rho_0,\sigma_0\}$ and $\mathcal E=\{\rho,\sigma\}$, this amounts to the
condition 
\[
P_e(\lambda,\rho_0,\sigma_0)\le P_e(\lambda,\rho,\sigma)+\epsilon,\qquad \forall
\lambda\in [0,1].
\]
Using Lemma \ref{lemma:qnm}, we see that this is equivalent to any of the inequalities
\begin{align*}
\|\rho_0-s\sigma_0\|_1&\ge \|\rho-s\sigma\|_1-2\epsilon(1+s),\qquad s\ge 0\\
\Tr[(\rho_0-s\sigma_0)_{+}]&\ge \Tr[(\rho-s\sigma)_{+}]-\epsilon(1+s),\qquad s\ge
0\\
\Tr[(\rho_0-s\sigma_0)_{-}]&\ge \Tr[(\rho-s\sigma)_{-}]-\epsilon(1+s), \qquad s\ge 0.
\end{align*}
It is easily seen that this is true if there is some channel $\Lambda$ such that
$\|\Lambda(\rho_0)-\rho\|_1\le\epsilon$ and $\|\Lambda(\sigma_0)-\sigma\|_1\le \epsilon$.
In the classical case the converse holds, but  in the quantum case this is not true. 
We can obtain some form of the converse statement if $\rho_0=\Phi(\rho)$ and
$\sigma_0=\Phi(\sigma)$, in a similar way as in Theorem \ref{thm:recoverability_e}.

\end{remark}

\section*{Acknowledgement}

I am grateful to Mark Wilde and Fumio Hiai for their comments on an earlier version of the
manuscript.
The research was supported by the grant VEGA 1/0142/20 and  the Slovak
Research and Development Agency grant APVV-20-0069.

\appendix

\section{Appendix: Proof of Theorem \ref{thm:petz}}

The implication (i) $\implies$ (ii) follows by the data processing inequality
\eqref{eq:dpi}, (ii) $\implies$ (iii) (with $t=0$) was proved in
\cite{jencova2006sufficiency, petz1988sufficiency}. The implications (iii) $\implies$ (i)
and (iv) $\implies$ (v) $\implies$ (i) are easy, so the only thing left to prove is (i)
$\implies$ (iv). 

By the assumption $D(\rho\|\sigma)<\infty$, we have that $\supp(\rho)\le \supp(\sigma)$
and also $\supp(\Phi(\rho))\le \supp(\Phi(\sigma)$. We may therefore assume that both
$\sigma$ and $\Phi(\sigma)$ are faithful, by restriction to the respective supports.

With this assumption, we will also need to recall some further properties of sufficient channels from
\cite{petz1988sufficiency,jencova2006sufficiency}. Let us denote by $u_t$ and $v_t$ the Connes cocycles
\[
u_t=\rho^{it}\sigma^{-it},\quad v_t=\Phi(\rho)^{it}\Phi(\sigma)^{-it},\qquad t\in \mathbb
R,
\]
then $u_t$ and $v_t$ are one-parameter families of isometries satisfying the conditions
\[
\sigma^{is}u_t\sigma^{-is}=u_s^*u_{t+s},\quad
\Phi(\sigma)^{is}v_t\Phi(\sigma)^{-is}=v_s^*v_{t+s},\qquad s,t\in \mathbb R.
\]
By the results of \cite{petz1988sufficiency,jencova2006sufficiency}, we can see that if
$\Phi$ is sufficient with respect to $\{\rho,\sigma\}$, then $u_t$ is in the
multiplicative domain of the unital completely positive map $\Phi_\sigma^*$ and similarly $v_t$ is in the multiplicative domain of
$\Phi^*$, see \cite[Theorem 3.18]{paulsen2002completely} for the definition and properties of
multiplicative domains. Moreover, $\Phi^*(v_t)=u_t$ and $\Phi_\sigma^*(u_t)=v_t$. It
follows that for any $s,t\in \mathbb R$,
\[
\Phi^*(\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it})=\Phi^*(v_t^*v_{s+t})=u_t^*u_{s+t}=\sigma^{it}u_s\sigma^{-it}
\]
and similarly
$\Phi_\sigma^*(\sigma^{it}u_s\sigma^{-it})=\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it}$. 
Let now $A\in B(\Ha)$, we have for all $s,t\in \mathbb R$,
\[
\<\sigma^{it}u_sA^*\sigma^{-it},\sigma^{it}u_s\sigma^{-it}\>_\sigma=\<u_sA^*,u_s\>_\sigma=\Tr[A\rho^{is}\sigma^{1/2-is}\rho^{is}_s\sigma^{1/2-is}]
\]
and the  analytic continuation to $s=-\tfrac12i$ of the last expression becomes
$\Tr[A\rho]$. On the other hand, using the above properties of the cocycles and of
multiplicative domains, we get by the definition of the Petz dual
\begin{align*}
\<\sigma^{it}u_sA^*\sigma^{-it},\sigma^{it}u_s\sigma^{-it}\>_\sigma&=\<\sigma^{it}u_sA^*\sigma^{-it},\Phi^*(\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it})\>_\sigma\\
&=
\<\Phi_\sigma^*(\sigma^{it}u_sA^*\sigma^{-it}),\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it}\>_{\Phi(\sigma)}\\
&=\<\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it}\Phi_\sigma^*(\sigma^{it}A^*\sigma^{-it}),\Phi(\sigma)^{it}v_s\Phi(\sigma)^{-it}\>_{\Phi(\sigma)}\\
&=\Tr[\Phi^*_{t,\sigma}(A)\Phi(\rho)^{is}\Phi(\sigma)^{1/2-is}\Phi(\rho)^{is}\Phi(\sigma)^{1/2-is}],
\end{align*}
here the analytic continuation to $s=-\tfrac12i$ equals 
\[
\Tr[\Phi^*_{t,\sigma}(A)\Phi(\rho)]=\Tr[A\Phi_{t,\sigma}(\Phi(\rho))].
\]
Since $A$ and $t$ were arbitrary, this finishes the proof of (i) $\implies$ (iv).


\end{document}